# Generic groups and the weak amalgamation property

### v2

@@ 1 Introduction
The recent paper [6] considers countable groups as elements of a Polish space in the following way. An enumerated group is the set $\omega$ together with a multiplication function $\cdot:\omega\times\omega\rightarrow\omega$, an inversion function :${}^{-1}:\omega\rightarrow\omega$ , and an identity element $e\in\omega$ defining a group. The space $\mathcal{G}$ of enumerated group is the closed subset of the space $\mathcal{X}=\omega^{\omega\times\omega}\times\omega^{\omega}\times\omega$ (under the product topology). Any abstract property of groups, say $P$, naturally defines the invariant subset $\mathcal{G}_{P}$ of $\mathcal{G}$. Assuming that $\mathcal{G}_{P}$ is Polish it is studied in [6] which group-theoretic properties are generic in $\mathcal{G}_{P}$. In particular, are there generic groups in $\mathcal{G}_{P}$? (A group $G$ is generic if all isomorphic copies of $G$ form a comeagre subset of $\mathcal{G}_{P}$.) The main results of that paper state that there is no generic group in $\mathcal{G}$ and moreover, when $P$ is the property of left orderability the space $\mathcal{G}_{P}$ does not have a generic group. The authors deduce this using involved arguments in the style of model-theoretic forcing.
A. Ivanov studied the problem of existence of generic objects in [7] in full generality. He introduced some general condition which is now called the weak amalgamation property (WAP) and showed that it together with the joint embedding property characterizes the existence of generics. 111 A. Ivanov used in [7] the name the almost amalgamation property. This condition was considered later in the case of automorphisms in the very influential paper [8] were the name WAP was used.
The goal of our paper is to demonstrate that the results of [6] mentioned above can be deduces by an application of [7]. In fact in Theorems 3.2 and 4.3 we formulate a very general property which implies the absence of generic objects. These meta-theorems can be applied in a number of cases. Along with ones studied in [6] we also prove:
@@
We also demonstrate how [7] works in the contrary direction, i.e. for proving of existence of generics in some natural group varieties. In particular in Section 2.3 we prove the following theorem.
Theorem 2.7. Let $c,p\in\omega\setminus\{0,1\}$ and $p$ be prime $>\mathsf{max}(2,c)$. Then the space of nilpotent groups of degree $c$ and of exponent $p$ is a closed subspace of $\mathcal{G}$ which has a generic group.
Furthermore, the approach of [6] is extended to actions of groups on first-order structures. In particular, we show that there is no generic group action on $(\mathbb{Q},<)$, see Corollary 3.5.
@@ 1.1 General preliminaries
Fix a countable atomic structure in a language $L$. Let $T$ be an expansion of $Th(M)$ in some $L\cup\bar{r}$ where $\bar{r}=(r_{1},\ldots,r_{\iota})$ is a sequence of additional relational/functional symbols. We assume that $T$ is axiomatizable by sentences of the following form:
$(\forall\bar{x})\left(\bigvee_{i}(\phi_{i}(\bar{x})\land\psi_{i}(\bar{x})\right),$
where $\phi_{i}(\bar{x})$, is a quantifier-free formula in the language $L\cup\bar{r}$, and $\psi_{i}(\bar{x})$ is a first order formula of the language $L$. Let $\mathcal{X}_{M,T}$ be the space of all $\bar{r}$-expansions of $M$ to models of $T$. This is a topological space with respect to so called logic topology. In the following few paragraphs we introduce some notions previously defined in [7]. They will be helpful in a description of the logic topology too.
@@
It is assumed in [7] that $Th(M)$ is $\omega$-categorical. It is worth mentioning that the arguments used in the proof of Theorem 1.2 in [7] work under the weaker assumption that $M$ is atomic. However in the present paper we will work under the assumption of $\omega$-categoricity.
For a finite $A\subset M$ we define a diagram of $\bar{r}$ on $A$ as follows. To every functional symbol $r_{i}$ from $\bar{r}$, say of arity $\ell$, a partial function $\mathsf{r}_{i}:A^{\ell}\to A$ is associated. When $r_{i}$ is a relational symbol we include into the diagram a single formula from every pair $\{\mathsf{r}_{i}(\bar{a}^{\prime}),\lnot\mathsf{r}_{i}(\bar{a}^{\prime})\}$, where $\bar{a}^{\prime}$ is a tuple from $A$ of the corresponding length. We will say that $A$ is the domain of this diagram. For ease of notation we will usually view $A$ as a tuple of pairwise distinct elements, say $\bar{a}$. Then given a diagram $D$ on $A$ we write $\bar{a}=\mathsf{Dom}(D)$. It will be also convenient to use the form $D(\bar{a})$, where the domain is already given in the expression.
Let $\mathcal{B}_{T}$ be the set of all theories such that each of them consists of a diagram $D(\bar{a})$, $\bar{a}\subset M$, the theory $T$ and the constant extension $Th(M,\bar{a})$ such that $D(\bar{a})\cup T\cup Th(M,\bar{a})$ is satisfied in some $(M,{\bf\bar{r}})\models T$ (here we view $D(\bar{a})$ as the atomic diagram of the corresponding partial structure). Since $D(\bar{a})$ uniquely defines the corresponding member of $\mathcal{B}_{T}$ we will identify this member with the diagram. Furthermore, since $M$ is atomic, each element of $\mathcal{B}_{T}$ is determined by a formula of the form $\phi(\bar{a})\wedge\psi(\bar{a})$, where $\psi(\bar{x})$ is a complete formula for $M$ realized by $\bar{a}$ and $\phi(\bar{a})$ is a quantifier-free formula in the language $L\cup\bar{r}$. The corresponding $\phi(\bar{x})\wedge\psi(\bar{x})$ is called basic.
@@
$[D]=\{(M,{\bf\bar{r}})\in\mathcal{X}_{M,T}\,|\,(M,{\bf\bar{r}})\mbox{ satisfies }D(\bar{a})\}$
is clopen in the logic topology. In fact the family $\{[D]\,|\,D(\bar{a})\in\mathcal{B}_{T}\}$ is usually taken as a base of this topology. It is metrizable. Indeed, take an enumeration of $M^{<\omega}$, say $M^{<\omega}=\{\bar{a}_{0},\bar{a}_{1},\ldots\}$. Define a metric $d$ metrizing $\mathcal{X}_{M,T}$ as follows:
@@
Let $D\in\mathcal{B}_{T}$ and $\mathsf{Dom}(D)=\bar{a}$. Then $D$ can be viewed as an expansion of $M$ by finite relations corresponding to $\bar{r}$. When $\mathsf{r}_{i}$ is a functional symbol the relation is $\mathsf{Graph}(\mathsf{r}_{i})$, the graph of the corresponding partial function on $\bar{a}$. For any automorphism $\alpha\in\mathsf{Aut}(M)$ we define $\alpha(D)$ to be the diagram obtained from $D(\bar{a})$ by substitution $\alpha(\bar{a})$ instead of $\bar{a}$.
@@
Since the space $\mathcal{X}_{M,T}$ is Polish, Bair category method can be applied.
@@
$\mathcal{X}_{M}$ has a generic expansion $(M,{\bf\bar{r}})$ if and only if the family $\mathcal{B}_{T}$ has JEP and WAP.
@@ 1.2 The case of groups
I. The basic case.
The structure $M$ is just $\omega$ and $L$ consists of one constant symbol $1$ which is interpreted by number $1$. Let $\bar{r}$ consist of the binary function of multiplication $\cdot$ and a unary function -1. Let $T$ be the universal theory of groups with the unit $1$  [^2] It is not necessary to fix $1$ in the language of $M$. For example, the paper [6] does not do this. We fix it just for convenience of notations in Scenario II below..
Each element of $\mathcal{B}_{T}$ consists of a tuple $\bar{c}\subset\omega$ containing $1$ and partial functions for $\cdot$ and -1 on $\bar{c}$. This is a finite partial group. The space $\mathcal{X}_{M,T}$ is the logic space of all countable groups, where the unit is fixed. We emphasize that this is exactly the space $\mathcal{G}$ of enumerated groups from [6]. Indeed, Corollary 3.1.3 of [6] gives a basis of clopen sets of $\mathcal{G}$ as follows. Take a finite system of group equations and inequations $\Sigma(\bar{x})$ and a tuple $\bar{a}\in\omega$ with $|\bar{x}|=|\bar{a}|$. Let $[\Sigma(\bar{a})]$ be the set of all group operations on $\omega$ for which the tuple $\bar{a}$ satisfies $\Sigma(\bar{x})$. Then the family of all sets of the form $[\Sigma(\bar{a})]$ is the basis of $\mathcal{G}$. On the other hand note that every element $D(\bar{c})\in\mathcal{B}_{T}$ determines a finite system, say $\Sigma_{D}(\bar{c})$, of group equations and inequations (for words which are defined) over $\bar{c}$ so that $[D]=[\Sigma_{D}(\bar{c})]$. Furthermore, each $[\Sigma(\bar{a})]$ is a union of basic clopen sets for some family of diagrams $D(\bar{c})\in\mathcal{B}_{T}$. Indeed, consider all finite partial multiplication tables on tuples $\bar{c}\subset\omega$ with $\bar{a}\subseteq\bar{c}$ such that every subword $u(\bar{a}^{\prime})$ of a word from $\Sigma(\bar{a})$ has a value in $\bar{c}$ and $\Sigma(\bar{a})$ is satisfied under this table. Each multiplication table of this form is viewed as a diagram $D(\bar{c})\in\mathcal{B}_{T}$.
@@
In the paragraph below we will have several sorts in $M$. The sort just described will always occur. It will be denoted by $\mathsf{Gp}$. In particular if $(M,{\bf\bar{r}})\in\mathcal{X}_{M,T}$ then we denote by $\mathsf{Gp_{\bar{r}}}$ the group defined by the corresponding operations from ${\bf\bar{r}}$ on this sort.
II. The case of an action.
Let $M_{0}$ be a countable atomic structure of some language $L_{0}$. Define $M$ to be $M_{0}$ with an additional sort $\omega$ called $\mathsf{Gp}$ and the constant symbol $1$ interpreted by $1$. The symbols $\bar{r}$ include $\cdot$, -1 and a new symbol $\mathsf{ac}$ for a function $M_{0}\times\mathsf{Gp}\rightarrow M_{0}$. Now the theory $T$ contains the universal axioms of groups on the sort $\mathsf{Gp}$ with the unit $1$. We also add the following axioms for an action.
@@
The space $\mathcal{X}_{M,T}$ is the logic space of all countable expansions of $M$ where the group structure is defined on $\mathsf{Gp}$, with a fixed unit $1$ and an action $\mathsf{ac}$
@@ 1.3 Generics are generic
In this section we indicate a connection of our approach (and approaches of [6] and [5]) with model-theoretic forcing. Let $\mathcal{X}$ be a Polish space with a countable basis of clopen sets $\mathcal{A}$. For $A\in\mathcal{A}$ and a Borel set $B$ we write $A\Vdash B$ if $A\setminus B$ is meagre in $A$ and $A\Vdash\sim B$ if $A\cap B$ is meagre in $A$. We say that $A$ decides $B$ if exactly one of these cases holds. If $\mathcal{B}$ is a family of Borel subsets, then following to Definition 2.2 of [11] we say that an element $x\in\mathcal{X}$ is $\mathcal{B}$-generic in $\mathcal{X}$ if for every $B\in\mathcal{B}$ there is a neighborhood $A\in\mathcal{A}$ of $x$ that decides $B$.
This notation is strongly connected with model-theoretic forcing. Recall that given a countable language $L$, a set of constant symbols $C$ and an $L$-theory $T$, a condition is a finite set $p$ of atomic formulas of $L\cup C$ without free variables and negations of such formulas such that $p$ is consistent with $T$. We will assume that $C=\omega$ and $T$ is universal. It is clear that each condition $p$ defines a set $[p]$ of all $L$-structures defined on $\omega$ which satisfy $p\cup T$. This is a clopen set in the logic space $\mathcal{X}_{T}$ of all models of $T$ defined on $\omega$. Let $\mathcal{A}$ be the basis of clopen set of $\mathcal{X}_{T}$ consisting of such $[p]$.
We now follow Section 8 of [9]. When $\phi$ is a formula over $L\cup C$ without free variables then Robinson’s forcing $p\Vdash\phi$ is defined as follows:
@@
Le $B=[\phi]$ be the Borel set of all $L$-structures on $\omega$ which satisfy $\phi\cup T$. Then $p\Vdash\phi$ implies $[p]\Vdash B$.
@@
Now consider the logic space $\mathcal{G}_{U}$ for some universal theory $U$. This is a Polish $Sym(\omega)$-space under the action by permutations on $\omega$. According to the previous paragraph if $M$ is a generic structure for $U$ then for every Borel set of the form $[\phi]$ there is a basic open $A$ of the form $[p]$ with $M\in[p]$ such that $[p]$ decides $[\phi]$. As we already know this property of $M$ is exactly $\mathcal{B}$-genericity in the terminology of [11] where in our case $\mathcal{B}$ is the family of all Borel subsets of $\mathcal{G}_{U}$ of the form $[\phi]$ for first-order $\phi$ 333the paper [11] also considers other examples of families $\mathcal{B}$. By Proposition 2.4 of [11] all $\mathcal{B}$-generic elements of $\mathcal{G}_{U}$ form a dense $G_{\delta}$-subset of $\mathcal{G}_{U}$. This proves the following statement.
@@
Since algebraic closedness coincides with existential closedness, this proposition strengthens Theorem 6.1 of [5]. Summarizing this section we conclude: if $G$ is a generic group of the space of enumerated groups $\mathcal{G}_{U}$, then $G$ is $U$-generic and existentially closed.

### v3

@@ 1 Introduction
The recent paper [8] considers countable groups as elements of a Polish space in the following way. An enumerated group is the set $\omega$ together with a multiplication function $\cdot:\omega\times\omega\rightarrow\omega$, an inversion function $^{-1}:\omega\rightarrow\omega$ , and an identity element $e\in\omega$ defining a group. The space $\mathcal{G}$ of enumerated groups is a closed subset of the space $\mathcal{X}=\omega^{\omega\times\omega}\times\omega^{\omega}\times\omega$ (under the product topology). Any abstract property of groups, say $P$, naturally defines the invariant subset $\mathcal{G}_{P}$ of $\mathcal{G}$. Assuming that $\mathcal{G}_{P}$ is Polish, it is studied in [8] which group-theoretic properties are generic in $\mathcal{G}_{P}$. In particular, are there generic groups in $\mathcal{G}_{P}$? (A group $G$ is generic if all isomorphic copies of $G$ form a comeagre subset of $\mathcal{G}_{P}$.) Among the rich variety of results of that paper we mention as the most notable the statement that there is no generic group in $\mathcal{G}$, and moreover, when $P$ is the property of left orderability, the space $\mathcal{G}_{P}$ does not have a generic group. The authors deduce this using involved arguments in the style of model-theoretic forcing.  [^1] Some of these results were independently (but later) obtained in [6]. This approach was also applied in [7].
A. Ivanov studied the problem of existence of generic objects in [9] in full generality. He introduced some general condition which is now called the weak amalgamation property (WAP) and showed that it together with the joint embedding property characterizes the existence of generics.  [^2] A. Ivanov used in [9] the name the almost amalgamation property. This condition was considered later in the case of automorphisms in the very influential paper [10] where the name WAP was used.
The goal of our paper is as follows. Using Ivanov’s approach we build a universal framework for investigations of generic objects in various classes of enumerated groups, semigroups and rings. Our meta-theorems are direct and short. In particular, see Theorem 3.2 which describes a very general property implying the absence of generic objects. The framework is applied in a number of cases. Firstly, we deduce the results of [8] mentioned above in a short fashion. Furthermore, we prove:
@@
We also apply our methods in the contrary direction, i.e. for proving of existence of generics in some natural group varieties. In particular, in Section 2.3 we prove the following theorem.
Theorem 2.9. Let $c,p\in\omega\setminus\{0,1\}$ and $p$ be prime $\mathsf{max}(2,c)$. Then the space of nilpotent groups of degree $c$ and of exponent $p$ is a closed subspace of $\mathcal{G}$ which has a generic group.
The framework also works in the case of actions of groups on first-order structures. In particular, we show that there is no generic group action on $(\mathbb{Q},<)$, see Corollary 3.7. This result is a contribution to the topic of generic representations of countable groups, see [5].
@@ 1.1 General preliminaries
Fix a countable atomic structure $M$ in a language $L$. Let $T$ be an expansion of $Th(M)$ in some $L\cup\bar{r}$ where $\bar{r}=(r_{1},\ldots,r_{\iota})$ is a sequence of additional relational/functional symbols. We assume that $T$ is axiomatizable by sentences of the following form:
$(\forall\bar{x})\left(\bigvee_{i}(\phi_{i}(\bar{x})\land\psi_{i}(\bar{x}))\right),$
where $\phi_{i}(\bar{x})$ is a quantifier-free formula of the language $L\cup\bar{r}$, and $\psi_{i}(\bar{x})$ is a first order formula of the language $L$ (in particular, $Th(M)\subseteq T$). Let $\mathcal{X}_{M,T}$ be the space of all $\bar{r}$-expansions of $M$ to models of $T$. This is a topological space with respect to so called logic topology. In the following few paragraphs we introduce some notions previously defined in [9]. They will also be helpful in a description of the logic topology.
@@
In [9] it is assumed that $Th(M)$ is $\omega$-categorical. It is worth mentioning that the arguments used in the proof of Theorem 1.2 in [9] work under the weaker assumption that $M$ is atomic. However, in the present paper we will work under the assumption of $\omega$-categoricity.
For a finite $A\subset M$ we define a diagram of $\bar{r}$ on $A$ as follows. To every functional symbol $r_{i}$ from $\bar{r}$, say of arity $\ell$, a partial function $\mathsf{r}_{i}:A^{\ell}\to A$ is associated. When $r_{i}$ is a relational symbol, we include into the diagram a single formula from every pair $\{\mathsf{r}_{i}(\bar{a}^{\prime}),\lnot\mathsf{r}_{i}(\bar{a}^{\prime})\}$, where $\bar{a}^{\prime}$ is a tuple from $A$ of the corresponding length. We will say that $A$ is the domain of this diagram. For ease of notation, we will usually view $A$ as a tuple of pairwise distinct elements, say $\bar{a}$. Then given a diagram $D$ on $A$, we write $\bar{a}=\mathsf{Dom}(D)$. It will also be convenient to use the form $D(\bar{a})$, where the domain is already given in the expression.
Given a diagram $D(\bar{a})$, $\bar{a}\subset M$, which is satisfied in some $(M,{\bf\bar{r}})\models T$, we view $D(\bar{a})$ as the atomic diagram of the corresponding partial substructure of $(M,{\bf\bar{r}})$.
Let $\mathcal{B}_{T}$ be the set of all theories of the form $D(\bar{a})\cup T\cup Th(M,\bar{a})$ that are satisfied in some expansions $(M,{\bf\bar{r}})\models T$. Since $D(\bar{a})$ uniquely defines the corresponding member of $\mathcal{B}_{T}$, we will identify this member with the diagram. Furthermore, since $M$ is atomic, each element of $\mathcal{B}_{T}$ is determined by a formula of the form $\phi(\bar{a})\wedge\psi(\bar{a})$, where $\psi(\bar{x})$ is a complete formula for $M$ realized by $\bar{a}$, and $\phi(\bar{a})$ is a quantifier-free formula in the language $L\cup\bar{r}$. The corresponding $\phi(\bar{x})\wedge\psi(\bar{x})$ is called basic.
@@
$D]=\{(M,{\bf\bar{r}})\in\mathcal{X}_{M,T}\,|\,(M,{\bf\bar{r}})\mbox{ satisfies }D(\bar{a})\}$
is clopen in the logic topology. In fact, the family $\{[D]\,|\,D(\bar{a})\in\mathcal{B}_{T}\}$ is usually taken as a basis of this topology. It is metrizable. Indeed, take an enumeration of $M^{<\omega}$, say $M^{<\omega}=\{\bar{a}_{0},\bar{a}_{1},\ldots\}$. Define a metric $d$ metrizing $\mathcal{X}_{M,T}$ as follows:
@@
Let $D\in\mathcal{B}_{T}$ and $\mathsf{Dom}(D)=\bar{a}$. Then $D$ can be viewed as an expansion of $M$ by finite relations corresponding to $\bar{r}$. When $\mathsf{r}_{i}$ is a functional symbol, the relation is $\mathsf{Graph}(\mathsf{r}_{i})$, the graph of the corresponding partial function on $\bar{a}$. For any automorphism $\alpha\in\mathsf{Aut}(M)$ we define $\alpha(D)$ to be the diagram obtained from $D(\bar{a})$ by substitution $\alpha(\bar{a})$ instead of $\bar{a}$.
@@
Since the space $\mathcal{X}_{M,T}$ is Polish, the Baire category method can be applied.
@@
$\mathcal{X}_{M,T}$ has a generic expansion $(M,{\bf\bar{r}})$ if and only if the family $\mathcal{B}_{T}$ has JEP and WAP.
@@ 1.2 The case of groups
I. The basic case
The structure $M$ is just $\omega$ and $L$ consists of one constant symbol $1$ which is interpreted by number $1$. Let $\bar{r}$ consist of the binary function of multiplication $\cdot$ and a unary function -1. Furthermore, a finite sequence of constants can be included into $\bar{r}$; this does not change the exposition. Let $T$ be the universal theory of groups with the unit $1$  [^3] It is not necessary to fix $1$ in the language of $M$. For example, the paper [8] does not do this. We fix it just for convenience of notations in Scenario II below..
Each element of $\mathcal{B}_{T}$ consists of a tuple $\bar{c}\subset\omega$ containing $1$ and partial functions for $\cdot$ and -1 on $\bar{c}$. This is a finite partial group. (When $\bar{r}$ has additional constants, it is assumed that distinguished constants are included into $\bar{c}$.) The space $\mathcal{X}_{M,T}$ is the logic space of all groups of size $\aleph_{0}$, where the unit is fixed. We emphasize that this is exactly the space $\mathcal{G}$ of enumerated groups from [8]. Indeed, Corollary 3.1.3 of [8] gives a basis of clopen sets of $\mathcal{G}$ as follows. Take a finite system of group equations and inequations $\Sigma(\bar{x})$ and a tuple $\bar{a}\subset\omega$ with $|\bar{x}|=|\bar{a}|$. Let $\Sigma(\bar{a})]$ be the set of all group operations on $\omega$ for which the tuple $\bar{a}$ satisfies $\Sigma(\bar{x})$. Then the family of all sets of the form $\Sigma(\bar{a})]$ is the basis of $\mathcal{G}$. On the other hand, note that every diagram $D(\bar{c})\in\mathcal{B}_{T}$ determines a finite system, say $\Sigma_{D}(\bar{c})$, of group equations and inequations over $\bar{c}$ (for words which values are defined under the corresponding restrictions of the group operations) so that $D]=[\Sigma_{D}(\bar{c})]$. Furthermore, each $\Sigma(\bar{a})]$ is a union of basic clopen sets determined by some family of diagrams $D(\bar{c})\in\mathcal{B}_{T}$. Indeed, consider all finite partial multiplication tables on tuples $\bar{c}\subset\omega$ with $\bar{a}\subseteq\bar{c}$ such that every subword $u(\bar{a}^{\prime})$ of a word from $\Sigma(\bar{a})$ has a value in $\bar{c}$ and $\Sigma(\bar{a})$ is satisfied under this table. Each multiplication table of this form is viewed as a diagram $D(\bar{c})\in\mathcal{B}_{T}$.
@@ I. The basic case
In the paragraph below we will have several sorts in $M$. The sort just described will always occur. It will be denoted by $\mathsf{Gp}$. The symbols $\bar{r}$ will always contain ones for multiplication and for inversion. In particular, if $(M,{\bf\bar{r}})\in\mathcal{X}_{M,T}$ then we denote by $\mathsf{Gp_{\bar{r}}}$ the group defined by the corresponding operations from $\bf\bar{r}$} on $\mathsf{Gp}$.
II. The case of an action
Let $M_{0}$ be a countable atomic structure of some language $L_{0}$. Define $M$ to be $M_{0}$ with an additional sort $\omega$ called $\mathsf{Gp}$ and the constant symbol $1$ interpreted by $1$. The symbols $\bar{r}$ include $\cdot$, -1 and a new symbol $\mathsf{ac}$ for a function $M_{0}\times\mathsf{Gp}\rightarrow M_{0}$. Now the theory $T$ contains the universal axioms of groups on the sort $\mathsf{Gp}$ with the unit $1$. We also add the following axioms for an action:
@@ II. The case of an action
The space $\mathcal{X}_{M,T}$ is the logic space of all countable expansions of $M$ where the group structure is defined on $\mathsf{Gp}$, with a fixed unit $1$ and an action $\mathsf{ac}$ .
@@ 1.3 Generics are generic
In this section we indicate a connection of our approach (and approaches of [8], [6] and [7]) with model-theoretic forcing. Let $\mathcal{X}$ be a Polish space with a countable basis of clopen sets $\mathcal{A}$. For $A\in\mathcal{A}$ and a Borel set $B$ we write $A\Vdash B$ if $A\setminus B$ is meagre in $A$, and $A\Vdash\sim B$ if $A\cap B$ is meagre in $A$. We say that $A$ decides $B$ if exactly one of these cases holds. If $\mathcal{B}$ is a family of Borel subsets, then following Definition 2.2 of [13], we say that an element $x\in\mathcal{X}$ is $\mathcal{B}$-generic in $\mathcal{X}$, if for every $B\in\mathcal{B}$ there is a neighborhood $A\in\mathcal{A}$ of $x$ that decides $B$.
This notation is strongly connected with model-theoretic forcing. Recall that given a countable language $L$, a set of constant symbols $C$ and an $L$-theory $T$, a condition is a finite set $p$ of $(L\cup C)$-formulas without free variables which are atomic or negations of atomic formulas, such that $p$ is consistent with $T$. We will assume that $C=\omega$ and $T$ is universal. It is clear that each condition $p$ defines a set $[p]$ of all $L$-structures defined on $\omega$ which satisfy $p\cup T$. This is a clopen set in the logic space $\mathcal{X}_{T}$ of all models of $T$ defined on $\omega$. Let $\mathcal{A}$ be the basis of clopen set of $\mathcal{X}_{T}$ consisting of such $[p]$.
We now follow Section 8 of [11]. When $\phi$ is a formula over $L\cup C$ without free variables, then Robinson’s forcing $p\Vdash\phi$ is defined as follows:
@@
Let $B=[\phi]$ be the Borel set of all $L$-structures on $\omega$ which satisfy $\phi\cup T$. Then $p\Vdash\phi$ implies $p]\Vdash B$.
@@
Now consider the logic space $\mathcal{G}_{U}$ for some universal theory $U$. This is a Polish $Sym(\omega)$-space under the action by permutations on $\omega$. According to the previous paragraph if $M$ is a generic structure for $U$ then for every Borel set of the form $\phi]$ there is a basic open $A$ of the form $[p]$ with $M\in[p]$ such that $[p]$ decides $\phi]$. As we already know, this property of $M$ is exactly $\mathcal{B}$-genericity in the terminology of [13], where in our case $\mathcal{B}$ is the family of all Borel subsets of $\mathcal{G}_{U}$ of the form $\phi]$ for first-order $\phi$ 444the paper [13] also considers other examples of families $\mathcal{B}$. By Proposition 2.4 of [13] all $\mathcal{B}$-generic elements of $\mathcal{G}_{U}$ form a dense $G_{\delta}$-subset of $\mathcal{G}_{U}$. This proves the following statement.
@@
Since algebraic closedness coincides with existential closedness, this proposition strengthens Lemma 5.2.7 of [8], Theorem 6.1 of [6] and Theorem 3.11 of [7]. Summarizing this section, we conclude: if $G$ is a generic group of the space of enumerated groups $\mathcal{G}_{U}$, then $G$ is $U$-generic and existentially closed.